OpenStudy (anonymous):

can someone please help me please with solving the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols.

OpenStudy (anonymous):

2x+7y=-7 -4x-3y=-19

OpenStudy (anonymous):

@iGreen hey buddy finally have some maths and im not banned anymore :D could you please help me?

OpenStudy (anonymous):

@danish071996 are you any good at maths?

OpenStudy (anonymous):

multiply eq 1 with 2

OpenStudy (anonymous):

it will become 4x+14=-14

OpenStudy (anonymous):

then aad this eq with-4x3y=-19

OpenStudy (anonymous):

so how would i do this part: Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols.

OpenStudy (anonymous):

find x and y from the eq i gave then put them in this form

OpenStudy (anonymous):

ok great thank you so much for your help !!!